OpenStudy (anonymous):

Determine the prime factorization of 90. A. 2^2• 3 • 5 B. 2^3• 11 C. 2^2• 11 D. 2 • 3^2• 5 Would the answer be C?

jimthompson5910 (jim_thompson5910):

2^2 is 4

jimthompson5910 (jim_thompson5910):

so 2^2*11 = 4*11 = 44 which is NOT 90

OpenStudy (anonymous):

what are the prime factors that multiply together to make 90?

OpenStudy (nurali):

90 = 2×3²×5

OpenStudy (anonymous):

keep dividing it continuously by possible lowest no and then you'll get!

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

Here's how you'd answer the question Start with 90 and divide by 2 to get 90/2 = 45. So 2 is a factor of 90 Then notice how 45/2 = 22.5 so 2 is NOT a factor of 45 ----------------- Now divide 45 by 3 to get 45/3 = 15. So 3 is a factor of 45 15 isn't a multiple of 2, but it's a multiple of 3. So 15/3 = 5 ------------------- We found the following prime numbers: 2, 3, 5 The '3' occurs twice, which is why the prime factorization is \(\Large 2*3^2*5\)

OpenStudy (anonymous):

ohh ok

jimthompson5910 (jim_thompson5910):

It might help to do a factorization tree like this 

OpenStudy (anonymous):

@jim_thompson5910 explained what i was trying to say!! it's the correct and best method

jimthompson5910 (jim_thompson5910):

circle the prime numbers in the factor tree 

jimthompson5910 (jim_thompson5910):

when you multiply all those circled prime numbers out, you get the final answer

OpenStudy (anonymous):

ohh ok, I didn't think of it that way, now i get it

jimthompson5910 (jim_thompson5910):

ok, that's good things are clicking now